OpenStudy (anonymous):

find the general solution.. 1. y''-2y'-2y=0 2. y''''-8y''-9y=0 3. y''''+13y''+36y=0 4. y'''''-15y''''+84y'''-220y''-275y'=0 5. y'''''+12y''''+104y'''+408y''-1156y'=0

myininaya (myininaya):

Have you tried solving that one equation I asked you to solve?

OpenStudy (anonymous):

no because its hard

OpenStudy (anonymous):

y''+2y'+2y=0 r^2 + 2r + 2 = 0 r^2 + 2r = - 2 r^2 + 2r + 1 = - 2 + 1 (r + 1)^2 = - 1 r + 1 = ± √(-1) r = - 1 ± i the general solution is: y = e^(-x)[c₁sin(x) + c₂cos(x)] y = c₁e^(-x)sin(x) + c₂e^(-x)cos(x)

myininaya (myininaya):

r^2-2r-2=0

myininaya (myininaya):

Just use the quadratic formula.

OpenStudy (anonymous):

from 2 to 5 cud u answer it?

OpenStudy (anonymous):

wat # that i will use quadratc fmla?

myininaya (myininaya):

Whenever you have something that can be written in the form ax^2+bx+c=0, you may use the quadratic formula. You also may find synthetic division useful.

OpenStudy (anonymous):

can u help me to solve this equation, if its ok 4 u....

myininaya (myininaya):

Which one are you stuck on?

OpenStudy (anonymous):

3,4,5...

myininaya (myininaya):

So looking at number 3, we need to solve \[y^{(4)}+13y^{(2)}+36y^{(0)}=0\] means we need to look at how to solve \[r^4+13r^2+36=0\] Let u=r^2 so we have \[u^2+13u+36=0\] Use the quadratic formula or factor.

OpenStudy (anonymous):

tnx how about #4 and # 5..just giv me a little bit of sol...and i'll be the one to continue to solv

myininaya (myininaya):

Let me see your attempt on 4.

myininaya (myininaya):

You should find one zero pretty easily.

OpenStudy (anonymous):

can u answer it because i nid to sleep and this is my retrice. it will submit 2mrror mrning...f it is k 4 u...tnx in advance.....

OpenStudy (anonymous):

myininaya?..pls

OpenStudy (anonymous):

tnx 4 everything uve shared to me....ur a genius as wat i obsvd